27 Sep 2013

We are over the peak harvest bump! Looking through the harvest data sheets its clear that we have harvested well over half of the heads that were in cg1 this year. Lydia and I are dissecting the heads that were used in the quantitative genetics experiment this year. Out of just over 200 we are more than halfway finished extracting the achenes, each head takes approximately half an hour. Prepping the new experimental plot is still in progress though most of the trees have been cleared out at this point. We’ve begun collecting little bluestem and Bouteloua seeds which we will broadcast in the new plot this fall. We’ve also discovered windy days are NOT the best days to do this, even a slight wind will carry away the little bluestem. We also completed the first demography recheck of the fall and there are many more to come.

11 September 2013

This week has been full of demography and clearing a plot to plant seeds from the crossing experiment that took place this summer. We’ve been busy lopping, chain sawing, weed whipping and painting stumps with roundup in order to make way for the incoming prairie. Demography is almost done, just one site left to hit up for our first sweep through.

This afternoon while Stuart was busy trapping gophers Lydia and I spent the afternoon harvesting heads from the crossing experiment, we were on our second to last head when we encountered a curious situation. We were at the correct row and position and just could not find the plant that we had visited so many times this summer. We did however find a small hole in the ground near where an Echinacea plant should be. We pondered what to do, then Lydia finally exclaimed “I’m going to go-for-it”! Then she put her face to the ground, peered in and shouted “I see a mesh bag!!” Then in a flurry she reached down inside the hole and pulled out an already harvested Echinacea head….

…the gophers must have read the protocol because they seemed to know to leave 3 centimeters worth of stem from the base of the head.

September 1-3, 2013

The first of the month started off with Phenology on only 4 plants…but now we are down to the last two! Head harvesting is underway and it looks like we will have plenty to harvest come Thursday. We are staying busy trying to keep up with repainting heads in the crossing experiment before they are ready to be harvested, but the grasshoppers are doing their best to sabotage our efforts, check out this guy caught in the act of eating off painted bracts!

Yesterday Lydia and I mapped out a potential location for a new common garden where the progeny from the crosses from this summer will be planted this fall. This morning we are off to do demography and survey, as well as harvest the heads at Hegg lake. We will harvest all plants in the experimental plot there as well as the pallida plants Dayvis observed pollenators on and the nearest angustifolia plants to determine if any of the progeny from those heads are a product of hybridization…we may see some interesting results!

21 August 2013

It’s getting lonelier and lonelier at the Hjelm house these days. Today is was just Lydia, Sarah, Kory, and I. The morning consisted of phenology and seedling refinds at SPP. Then after lunch Lydia worked on her aphid addition/exclusion experiment and it seems that the aphids are disliking this heat wave almost as much as us, if not more, as she found an abundance of aphid skeletons this afternoon. I pollentated and repainted some heads while Kory and Sarah knocked off a few rows each of refinds in the cg1. Then the three of us hit up LC and found some seedlings. Here is a seedling from 09, as you probably could have surmised yourself.622640_10200336786414398_1112130388_o.jpg

The afternoon wrapped up with the most pressing of tasks…some unfinished business if you may. We began the process of finishing up some ice cream that has been in the freezer all summer. Hopefully it wont take too long to finish off the rest of the ice cream, but I’m sure we will manage somehow, working long hours if we must.

13 August 2013

We continued with phenology, measuring cg1, seedling refinds and crossing today. The day ended with a wonderful dinner put on by Stuart and Gretel et al., and not one but TWO bonfires at the farm house complete with s’more fixings. On Dayvis’s last night here he had his first s’more ever! Marie left in the evening and Dayvis will leave in the morning, our numbers are definitely dwindling but there is still plenty left to be done.

10 August 2013

A low key day here in Kensington, not much Echinacea Project work going on today. We bid our sad farewells (those of us who stuck around this weekend) to Sara Z as she left to head home and begin preparing for another year of teaching in Chicago. Today was filled with some cleaning, laundry and grocery shopping as per the usual; however, I did come across this curious sign today on my run, a historic site with arrows pointing in BOTH directions….


So curiosity struck and I decided to follow the arrow that pointed right, and I came across this little gem…

Thumbnail image for 173.JPG

A monument erected in honor of the towns founders. To find out what is in the direction of the historic site sign’s arrow you will have to go see for yourself.

7 August 2013

Another day of measuring, phenology, and crossing has come and gone. Phenology this morning was quite wet, afterwards some of us crossed and the rest worked on independent projects. Crossing has gotten quite challenging in the sense that we have a limited supply of pollen to work with as most of the sires are done flowering. The challenge is to get an even distribution of each sire (pollen donor) to the 4 dams (pollen acceptor) that it goes to, this has proven to be our biggest obstacle. And to top it off some dams haven’t even started to flower yet!

At lunch we got to see Dayvis’s preliminary power point for the presentation he will be giving on his research at the Chicago Botanic Garden next week; Marie will also be presenting a poster on her hybrid project there.

Then it was back to work with measuring the common garden and finishing up crossing jobs from this morning for the rest of the afternoon.backtowork.JPG

28 July 2013

Today Reina, Mike and I went out to do phenology this morning, along with Gretel, Stuart and Per it went by pretty quick! I found this confused ray florette this morning.

And below is a head with just one lone anther…enjoying its last day of flowering.

In the afternoon Gretel and I worked on continuing crossing in the qGen experiment. All in all it was a relatively quiet day here in Kensington.

25 July 2013

This morning started out with some of the crew at varying remnants working on independent projects and some were at the Hjelm house: Reina and Pam were working on their trichome data, while Sara Z worked on pinning ants, Lydia and I began some crosses but were soon rained out. Before we got rained out however I found this large fly larva in the head of the plant I was crossing…normally they are transparent-ish…this one however had a hay day with the pollen and as you can see is quite orange!full_magot.jpg

Once the rain put a stop to the crosses for the morning we worked on organizing and taking inventory of the pollen that has been collected thus far for the crossing experiment (qGen2).

After lunch today Stuart gave the crew a brief introduction into the wonderful world of R…OHHHHH the possibilities….

Then, later in the afternoon, we had many hands on deck helping with crosses that had been prioritized for the day. The most important thing about crossing is that you accessorize well while doing it, and what better way to accessorize than to have a handy bract painting bracelet! Below I am showing proper bracelet wearing technique.

These bracelets are all the rage right now at the Hjelm house, and they certainly are not going out of “style” anytime soon! (Or at least not until the crossing experiment is finished). Well, I must say we are very fashion forward here in Kensington, and I’m sure no one will be caught without one tomorrow because of the exciting happenings that are to be taking place tomorrow morning….but I shouldn’t say anymore just yet….stay tuned!

17 July 2013

It was another hot day on the prairie front. We all broke off into little groups today to take care of business. This morning Reina, Mike, and Pam assed herbivory in INB1 and INB2. Kory, Marie, Sarah B, and Dayvis were all scattered about working on their independent projects. Lydia and Sara Z assed the maternal plants in the common garden that are to be involved in the crosses of the quantitative genetic experiment, and on the other side of things Gretel, Shelley (Gretel’s Mother), Ruth and I went to Landfill Core to gather pollen from the sires that are to be involved in the crossing experiment. We used toothpicks to scrape off the pollen into vials that will be used for pollenating the maternal plants in the common garden. We were able to get pollen from about 40 of 55 plants at the site. The 15 others were just a little too immature to collect pollen from today.


This afternoon encompassed an hour of stipa searching for some and then continuation of some independent work on projects. Herbivory assessment continued as well. Stuart, Gretel, Ruth and I went for a nice long walk around SPP. We walked a U-shaped transect and used a randomized scheme for assigning sires to be used in the quantitative genetics experiment. We caged/bagged heads that were to be used and gps-ed the plants.

What a busy day!